OpenStudy (anonymous):

Raven’s parents bought a 20’ ski boat for $22,000. After 2 years, the boat was worth 17,500. Using x as the number of years and y as the boat value, graph the two points. Part 1: [2 points] Describe the slope of the line which connects the two points. Part 2: [3 points] How can your graph be used to predict the value of the boat in 5 years? Part 3: [2 points] What is the projected boat value in 5 years?

OpenStudy (anonymous):

please help me

OpenStudy (luigi0210):

I would use the slope formula..

OpenStudy (anonymous):

idk how to do any of this tho

OpenStudy (luigi0210):

\[\frac{22000-17500}{2-0}=\frac{4500}{2}=2250\] That is the slope of the line: y=22000-2250x

OpenStudy (luigi0210):

Part 1: The slope decreases with each year

OpenStudy (luigi0210):

For part 2&3, just plug in 5 into the equation

OpenStudy (anonymous):

huh?

OpenStudy (luigi0210):

^_^' I am awful at explaining.. sorry Plus I gtg. @Loser66

OpenStudy (anonymous):

crap

OpenStudy (anonymous):

ive been taking this exam for an hour and a half mainly cause of this question and i cant skip it ugh